OpenStudy (anonymous):

Differentiate F(x) = (x^4 + 3x^2 - 2)^5

OpenStudy (ash2326):

We have \[f(x)= (x^4+3x^2-2)^5\] Let's differentiate this with respect to x \[f'(x)=\frac{d}{dx}( x^4+3x^2-2)^5\] we get now \[f'(x)= 5(x^4+3x^2-2)^{(5-1)} \times \frac{d}{dx}(x^4+3x^2-2)\] we get \[f'(x)= 5(x^4+3x^2-2)^4 \times(4x^3+6x)\]